OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = −1.

OpenStudy (carolinar7):

Ok There is a formula for that

OpenStudy (carolinar7):

It is \[\sqrt{(x1-x2)^2-(y1-y2)^2}=\sqrt{y1-y2}\]

OpenStudy (carolinar7):

So plugin in the values \[\sqrt{(x-0)^2-(y-1)}=\sqrt{(y+1)^2}\]

OpenStudy (carolinar7):

The parenthesis will cancel out

OpenStudy (carolinar7):

you will be left with\[(x-0)^2-(y-1)^2=(y+1)^2\]

OpenStudy (carolinar7):

Leave the x values a lone

OpenStudy (carolinar7):

and put all the y values on the right side

OpenStudy (anonymous):

Ok, Im confused. What would the final solution come out to?